A source of light is at a temperature of 1000 °C. Its temperature is then increased to 2000 °C, Increasing the temperature causes the peak intensity wavelength to (a) decrease but not quite halve (b) more than double (c) halve (d) double (e) remain the same
Answer #1

By wein's displacement law we have

\lambdamax *T = 2898 um

\lambdamax = 2898/T

where T is in kelvin

\lambdamax = 2898/1000+273 = 2.27 um

\lambdamax = 2898 / 2000+273 = 1.27 um

when temperature is increases

so it will decrease but not quite halve

so option a) is the correct answer